OpenStudy (anonymous):

Use the quadratic formula to solve 16x2 + 96x + 128 = 0

OpenStudy (anonymous):

-4,-2

OpenStudy (anonymous):

x^2+6x+8=0 divide it entirely by 16 x^2+4x+2x+8=0 (x+4)(x+2)=0 x=-4 or -2

OpenStudy (anonymous):

thank you for the help